OpenStudy (anonymous):

What is the quotient of (t+3)/(2t) divided by (2t+6)/(t-4)?

OpenStudy (blurbendy):

It looks like this? \[\frac{ t + 3 }{ 2t } / \frac{ 2t -6 }{ t-4 }\]

OpenStudy (blurbendy):

/ means "divide"

OpenStudy (anonymous):

Yeah!

OpenStudy (anonymous):

@blurbendy I set it up with the right hand side as the inverse so that I could multiply across, and I was thinking that I solved it down properly, but my answer didn't match any of the answer choices.

OpenStudy (blurbendy):

okay, so you did \[\frac{ t + 3 }{ 2t } * \frac{ t-4 }{ t2 - 6} ?\]

OpenStudy (anonymous):

Yes, and I got (t+3)(t-4)/2t(t-4).

OpenStudy (anonymous):

So I wrote it out as (t^2-t-12)/(2t^2-8t)...

OpenStudy (blurbendy):

bottom should be 2t(2t +6)

OpenStudy (blurbendy):

yeah, i just saw that :)

OpenStudy (blurbendy):

thanks chief

OpenStudy (blurbendy):

do you see why the bottom is 2t(2t+6) @tasteless-teens ?

OpenStudy (anonymous):

Can you see this attachment?

OpenStudy (blurbendy):

yes, so you factored the top right when you got: (t+3)(t-4) the bottom is 2t(2t + 6) = 4t^2 + 12t = 4t(t + 3) = (t + 3) (t-4) / 4t(t+3) Do you see anything that can cancel out?

OpenStudy (anonymous):

the t+3

OpenStudy (blurbendy):

yes, so what does your answer become

OpenStudy (anonymous):

t-4/4t

OpenStudy (blurbendy):

good job!

OpenStudy (anonymous):

Wow, I really confused myself on that. Thank you! XD

OpenStudy (blurbendy):

yw

OpenStudy (anonymous):

Wait, so t does not equal 0 and 4, or t does not equal -3, 0, and 4?

OpenStudy (blurbendy):

look at your original equation. if t = 0, then 2(0) = 0 in the denominator which is undefined if t = 4, then 4 -4 in the denominator of the second expression results in 0 which is also undefined.

OpenStudy (anonymous):

So when we set it as 0, t does not equal 4?

OpenStudy (anonymous):

And that -3 was irrelevant?

OpenStudy (blurbendy):

no, it's saying t can't equal 0 or 4 because if you look at your original equation if you let t = 0, then for (t + 3/2t) (0 + 3) / 2(0) = 3/0 which is undefined If t = 4, then for (2t + 6) / (t- 4) = 8 +6 / 0 = 14/0 which is undefined

OpenStudy (blurbendy):

you never want 0 in the denominator, that's bad. the answer is the second one, so the one where t = -3 is wrong

OpenStudy (anonymous):

I just thought it was strange that the t =-3 was there. Thanks for pointing this out for me.

OpenStudy (blurbendy):

np